the probability of rolling a pair of dice that sums 7 is pie= 1/6

a) out of 4 independent rolls of the dice what is the mean number of timed the dice will show a 7?

b) out of 4 independent rolls of the dice what is the probability of rolling a 7 three or more timed ?
Answer #1

= 1/6

Number of trials, n = 4

P(X) = nCx x (1 - )n-x

a) Mean number of times the dice will show a seven = n

= 4 x 1/6

= 0.6667

b) P(rolling a 7 three or more times) = P(X = 3) + P(X = 4)

= 4C3x(1/6)3x(5/6) + (1/6)4

= 0.0162